OpenStudy (blackops2luvr):

A cardboard popcorn container is shaped like a cone that is 11 inches high and has a radius of 2.5 inches. What is the volume of the container? Use 3.14 for pi. Round your answer to the nearest hundredth.

OpenStudy (blackops2luvr):

@iGreen

OpenStudy (igreen):

Formula: \(\dfrac{1}{3} \pi r^2 h\) Plug in what we know: \(\dfrac{1}{3} (3.14) (2.5)^2 (11)\) Simplify exponent: \(\dfrac{1}{3} (3.14) (6.25) (11)\) Multiply: \((1.05) (6.25) (11)\) Can you multiply that? @blackops2luvr

OpenStudy (blackops2luvr):

yas.

OpenStudy (blackops2luvr):

the answer is D,.

OpenStudy (blackops2luvr):

@igreen

OpenStudy (igreen):

You didn't give options..lol.

OpenStudy (blackops2luvr):

27.53 in^3 40.35 in^3 46.07 in^3 71.96 in^3

OpenStudy (igreen):

Yep.